OpenStudy (anonymous):

Rachael's brother is 33 years old. He is 3 times what Rachael's age was 16 years ago. How old is Rachael?

OpenStudy (mathstudent55):

Call Rachael's age now r, and the brother's age now b. We are told b = 33 Since Rachael now is r years old, how old was she 16 years ago?

OpenStudy (anonymous):

33-16=17 17*3=51 51+16 Therefore she is 67 ? ... Not sure

OpenStudy (mathstudent55):

Rachael's age is r now. 16 years ago she was r - 16. The brother is now 3 times older than she was 16 years ago b = 3(r - 16), but we know b = 33, so 33 = 3(r - 16) 11 = r - 16 r = 27 Rachael is now 27 years old. Check: If Rachael is now 27 years old, 16 years ago she was 27 - 16 = 11 years old. 3 * 11 = 33 which is her brother's age now.

OpenStudy (anonymous):

oh ok thanks a lot @mathstudent55

OpenStudy (mathstudent55):

You're welcome.